OpenStudy (anonymous):

An origami expert took a square piece of paper measuring 8x8 inches, and folded it in half, then in half again, and again... making five folds altogether. How many square inches are there in the area of the final shape? An origami expert took a square piece of paper measuring 8x8 inches, and folded it in half, then in half again, and again... making five folds altogether. How many square inches are there in the area of the final shape? @Mathematics

OpenStudy (anonymous):

dont know

OpenStudy (saifoo.khan):

64/2 = 32 32/2 = 16 16/2 = 8 8/2 = 4 4/2 = 2

OpenStudy (saifoo.khan):

So 2 sqin.